# Optimizing compilation of error correction codes for $2\times N$ quantum dot arrays and its NP-hardness

### v1

@@ VI Optimal Compilation of regular LDPC codes
First we notice how a stabilizer tableau or parity check matrix maps to the input $\mathcal{S}$ of our ancillary qubit re-indexing problem. For example, take the Steane code:
@@
We now notice that if our error correction code has constant column weight $w_{c}$ (each column has the same number of non-zero entries), then this would mean that all elements that occur in $\mathcal{S}$, occur the same number of times. Because each element occurs $w_{c}$ times, we can then make $w_{c}$ identical chains of the form $n-1,n-2,...,0$, each of which has no gaps, and will pack perfectly. Therefore, all LDPC codes with constant column weight can be compiled to run on $2\times n$ architectures in exactly $w_{c}$ shuttles. It is worth noting that this would then include all “regular” LDPC codes, as they have constant column and row weight [19]. Furthermore, the $[[144,12,12]]$ code used in IBM’s recent 2024 paper [18] also has constant column weight and would be an excellent choice of error correcting code for this architecture.
VI.1 Optimal Compilation on Variable Column Weight Codes
@@
VI.2 Structured vs. Unstructured Codes
As discussed earlier in this section, any code with constant column weight will compile optimally under Shor-style syndrome extraction. In contrast, for naive syndrome circuits, codes with rich structure to them such as bivariate bicycle (BB) codes (including IBM’s “gross” code [18]) and toric codes, tend to occasionally perform worse after ancillary qubit re-indexing, while their Shor syndrome circuit compilation often saturates the lower bound on possible required shuttles. The reason for this poor performance on naive syndrome circuits is that the structure of these codes happens to naturally lead to quite decent compilation only using gate shuffling. However, the heuristics proposed here for naive circuits are simply that, heuristics. These methods based on sorting by some heuristic metric cannot capture the richness of the code’s original structure, and therefore applying them to permute the ancillary indices can increase the number of required shuttles by dissipating some of the original code structure. We do note that on the already quite well-structured codes we tested, the gate shuffling solution alone finds good, possibly even optimal, solutions. The naive syndrome ancillary re-indexing heuristics can work well on some codes, but they trying to solve a much harder problem than Shor syndrome re-indexing, which is already NP-hard as proven in Appendix .2. Furthermore, we reiterate that by using Shor syndrome extraction instead, we can do our best to solve a single NP-hard problem, instead of multiple in parallel (which is the naive case), and that codes with constant weight have a trivial optimal solution under Shor syndrome re-indexing.
In contrast to codes with good structure, in this section we present results on randomly generated codes for the purpose of examining the performance of our compilation techniques when codes do not have such an exploitable structure.
First, we consider performance of classical random regular LDPC codes under compilation. All of these will compile to $w_{c}$ shuttles under Shor-style syndrome extraction, where $w_{c}$ is the column weight, so we instead focus on performance of the naive circuits, which can be seen in Fig. 10. We observe that AHR for naive syndrome circuits performs better as the row weight increases, asymptotically approaching $n$. As expected, the increasing the row weight has no effect on the compilation performance of Shor syndrome circuits. Also as expected, increasing the column weight causes the number of shuttles required for Shor syndrome circuits to increase accordingly, again confirming that for regular codes, compilation into $w_{c}$ is trivially obtainable for Shor syndrome circuits. Furthermore, increasing the column weight also causes the number of shuttles found after naive circuit compilation to increase rather significantly, although seemingly linearly.
Figure 10: Compilation performance of random classical regular LDPC codes, comparing naive and Shor syndrome circuits and their compilation. Each data point represents the average of the compilation performance of 20 randomly generated regular LDPC codes. In a) we hold the column weight constant, while the row weight of the code is varied, while in b) the row weight is held constant while the column weight is varied.
Code Uncompiled Gate Shuffled AHR SSSC Allowing Spaces (Overhead)
[[100, 10, ?]] w/ 20 layers $445\scriptstyle\pm 20$ $443\scriptstyle\pm 19$ $144\scriptstyle\pm 11$ $82\scriptstyle\pm 7$ $25\scriptstyle\pm 1$ ($390\scriptstyle\pm 60$ blank dots)
[[100, 20, ?]] w/ 20 layers $399\scriptstyle\pm 20$ $398\scriptstyle\pm 20$ $132\scriptstyle\pm 12$ $76\scriptstyle\pm 7$ $22\scriptstyle\pm 1$ ($350\scriptstyle\pm 48$ blank dots)
[[200, 10, ?]] w/ 20 layers $975\scriptstyle\pm 32$ $972\scriptstyle\pm 32$ $291\scriptstyle\pm 17$ $160\scriptstyle\pm 10$ $27\scriptstyle\pm 1$ ($1125\scriptstyle\pm 164$ blank dots)
[[200, 20, ?]] w/ 20 layers $920\scriptstyle\pm 30$ $918\scriptstyle\pm 30$ $279\scriptstyle\pm 20$ $153\scriptstyle\pm 11$ $25\scriptstyle\pm 1$ ($1090\scriptstyle\pm 158$ blank dots)
Table 4: Shor syndrome compilation data for random quantum brickwork codes [20] Results displayed here were collected and averaged over 50 random circuits. The same compilation techniques were used as in Table 2, except now there is an additional column displaying the absolute lower bound on the number of shuttles that could be required, which is achieved by allowing blanks in the ancilla array. The overhead number refers to an upper bound on the number of blank quantum dots that would need to be inserted into the ancilla array to achieve this, as discussed in Section VI.1. This is an upper bound because we used the most naive method of filling all gaps in chains with blank ancilla, and we did not investigate more advanced algorithms for optimizing this in this work. In other words the number of required blanks in the ancillary array reported here is simply equal to the sum of the number of gaps over each chain in the minimal set of chains.
Furthermore, we evaluate our techniques for Shor syndrome circuit compilation on codes generated from random quantum circuits, so called “brickwork” codes [20]. Referring to Table 4, we see that our techniques consistently lower the the number of required shuttles for extracting syndrome on these random codes.
@@ VII Discussion and Conclusion
We have demonstrated methods for scheduling the shuttles of qubits in $2\times n$ one-dimensional architectures, highlighting a family of LDPC codes that under Shor-syndrome extraction only ever requires a number of shuttles exactly equal to the code’s column weight, which notably includes bivariate bicycle quantum LDPC codes, which were shown in recent work to have a threshold on par with the surface code [18].
@@
Furthermore, we show quite good results at reducing the shuttle requirements of even quite unstructured quantum codes (see Table 4 and Fig. 10). With the exception of the LDPC and random codes, for all codes compiled, we also performed comprehensive gate-by-gate simulation of the error correction pipeline in order to demonstrate that ignoring the commutativity of gates usually still results in circuits that function properly. However, we show in Table 3 fault tolerant compilation numbers of running the X and Z circuits separately, thereby guaranteeing commutativity, as we can perform the X checks while shuttling in one direction and then the Z checks while shuttling in the other, as proposed in [5]. For all codes tested, the sum of shuttles from compiling X and Z separately is roughly equal to doing them together. However, because the compiled $X$ and $Z$ circuits might require different labeling of ancillary qubits, this might further complicate preparation of the required cat states needed for this form of fault-tolerant syndrome extraction, as multiple independent cat states need to be interwoven into a 1D array of qubits (a possible solution to this is discussed in Appendix .3). For future work, we plan to investigate the design of more elegant algorithms for compilation, perhaps exploring the trade-off between minimizing the size of the ancilla array and minimizing the number of shuttles. We will also consider other paradigms and pieces of the quantum error correction pipeline possible in the specific hardware implementation; for instance, incorporating flag qubits, architectures with more than two rails, how to perform transversal operations, and discovery of other family codes that can have their syndrome extracted in ways that require only a few shuttling operations.
@@
Nevertheless, what we have presented here is already a bombastic achievement towards embarking on what we believe to be a rich area of research, both from the perspectives of theoretical computer science as well as practical quantum error correction. That is, the rather novel algorithmic questions raised by optimizing error correcting codes to run on practical hardware are challenging enough to be interesting, yet seemingly constrained enough to allow for progress to be made. The culmination of which is the practical result of having more efficient quantum error correction schemes specifically tailored to a specific and promising type of quantum hardware. We hope to see this work expanded upon, as well as other works that explore these sorts of optimizations for other qubit platforms as we continue to approach a future of fault-tolerant quantum computation.
@@ Acknowledgements.
e would like to thank our colleagues Filip Rozepedek and Guus Avis for their fruitful discussions regarding gate commutativity. We also thank Nithin Raveendran for his recommendation to see how bivariate bicycle codes compile on this architecture. We would like to acknowledge colleagues Ferdinand Kuemmeth and Fabrizio Berritta as well. This work was supported by funding from the Inge Lehmann Programme of the Independent Research Fund Denmark, the US Army Research Office (ARO) under Award No. W911NF-24-2-0043, and the NSF under grants 1941583, 2346089, and 2402861.
@@
In this appendix, we provide examples of the block stacking visualization on valid error correcting codes, an illustration of how to translate between the “ladder” and “staircase” notations, a proof for the NP hardness of the ancillary qubit re-indexing problem, a discussion of how to prepare the necessary interwoven cat states needed for compiled Shor syndrome circuits, and a variety of plots demonstrating how our compilation affects error rates of various codes.
@@
Refer to Figures 11 and 12 to see how the simple heuristics from Section III can be applied to the naive syndrome extracting circuits of Steane’s [[7,1,3]] code [14] and Shor’s [[9,1,3]] code [21]. In Section IV, we introduced a condensed version of the staircase notation that is specifically for Shor syndrome circuits. Fig. 13 illustrates how these two notations are equivalent.
Figure 11: Application of heuristics on the Steane code. Here we have the primitive sets of the naive syndrome extraction circuit of the Steane code expressed in block form. We can see that either of the first two heuristics discussed in Section III will allow use to find an optimal solution for this case of ancillary qubit re-indexing.
Figure 12: Application of heuristics on the Shor code. Here we have the primitive sets of the naive syndrome extraction circuit of the Shor code expressed in block form. We can see that either of the first two heuristics discussed in Section III will allow use to find an optimal solution for this case of ancillary qubit re-indexing.
@@
For this reduction, we take an arbitrary 3 partition problem and map it to our Shor re-indexing problem by mapping the elements of $A$ to a multiset $\mathcal{S}$, where $\mathcal{S}$ is the input to our problem as defined above. This $\mathcal{S}$ can then be decomposed into no fewer than $13m+$ 1 sets, each corresponding to a chain $c_{i}$, where $i\in\{0,1,...,3m\}$. Rather than specifying each element in $\mathcal{S}$, we instead define each $c_{i}$ below, noting that $\mathcal{S}=\sum_{i=0}^{3m}c_{i}$ (addition of multisets here is defined as combining them, e.g. $\{1,2\}+\{2,3\}=\{1,2,2,3\}$.) Later we prove that with how our $\mathcal{S}$ is defined, no other set of chains would be feasible in obtaining a solution to the reindexing problem, thus excluding the possibility of a solution to Shor re-indexing existing that does not also map to a valid solution of the given 3-partition problem. We define $c_{0},c_{1},...,c_{3m}$ as follows:
@@
Finally we notice that 00 occurs $3m$ times, so this construction needs at least $3m$ shuttles. Furthermore, due to 3) above, our construction will require a minimum of $13m+$ 1 shuttles, and our problem is to determine whether or not a bijection $\pi:[s]\rightarrow[s]$, recall $s=|\mathcal{S}|=\sum_{i=0}^{3m}|c_{i}|$, exists such that:
@@
To structure our proof, we first prove that if a solution exists in the 3-partition instance, then a solution of exactly $13m+$ 1 shuttles must exist for our problem. We prove afterwards that if a solution to our problem exists, then a solution to the 3-partition problem must also exist.
@@
Therefore each triple of chains would be packable into the gaps that occur within $c_{0}$. Therefore, all chains can be packed into $s$ positions exactly, implying the existence of a $13m+$ 1 shuttle solution.
@@
Therefore $m(T+4a^{*})-1$ is the number of spaces between the first and last element, where the $-1$ is due to the number of spaces between, for example, $2$ and 00 being $1$ in this context. As provided by Lemma 2, in those spaces, we have $(m-1)$ ridges of size $a^{*}$, one ridge of size $a^{*}-1$. So there are a total of $(m-1)a^{*}+(a^{*}-1)$ elements in this spaces, meaning that the total size of all gaps in $c_{0}$ is given equal to :
@@
Direction 2: Shor Re-indexing $\rightarrow$ 3-Partition: Now we prove that any $13m+$ 1 shuttle solution to Shor re-indexing would also be a solution to the 3-partition problem. We prove this by proving that our construction could only ever provide a $13m+$ 1 shuttle solution if our definitions of $c_{0},c_{1},...,c_{3m}$ are followed exactly, which would then mean that each chain, aside from $c_{0}$, represents an element from $A$ encoded in unary, and as discussed in Direction 1, the packing of those chains corresponds to triples from $A$ each summing to $T$. In other words, we designed our reduction so that technically speaking, the idea of packing chains is not enforced, however by design, the only possible $13m+$ 1 shuttle solution will also be the solution to packing our chains as defined, and no other collection of chains would be feasible. To show that no other collection if feasible, we first prove the remaining two lemmas.
@@
We observe that $a^{*}<\frac{T}{4}$, and that $a<\frac{T}{2}$ for all $a\in A$. Therefore $\max_{e\in\sum_{i=1}^{3m}c_{i}}e<\frac{T}{4}+\frac{T}{2}$. The smallest value in $c_{0}$ is $|S|+T$, so if the largest element in $\sum_{i=1}^{3m}c_{i}$ and the smallest element in $c_{0}$ were in the same chain, a gap of size at least $|S|+T/4-1$ would be created. Because we are looking for a bijection $\pi$ that maps from $1,...,|S|$ to $1,...,|S|$, the longest possible chain (including its gaps) would be limited to length $|S|$, and the largest gap possible would be $|S|-2$. Therefore, because $|S|+T/4-1>|S|-2$, as $T/4>0$, it would not be possible for any element of $c_{o}$ to participate in one of the other chains. ∎
@@
We use the above lemmas to prove that no other collection of chains would result in a $13m+$ 1 solution. For the sake of contradiction, let’s assume that such an alternate set exists. By Lemma 3, we know that $c_{0}$ must remain unchanged, as $c_{0}$ constitutes $1$ shuttle, and the other $3m$ shuttles must come from the other elements, as 00 occurs $3m$ times. Therefore this alternate set of chains would be $\{c_{0},c_{1}^{\prime},...,c_{3m}^{\prime}\}$, and a gap must occur at least once in a chain in $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ by Lemma 4. By Lemma 1, we know there are $m$ gaps of size $T+3a^{*}$, and thus to have a $13m+$ 1 solution, we must be able to pack $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ into those $m$ gaps of size $T+3a^{*}$. Therefore the gaps introduced by using an alternate set of chains must be filled by other chains. As all gaps are by definition of size less than $a^{*}$, the ridges (Lemma 2) within $c_{0}$ cannot be used to fill them, which then leaves the only possibility to be that within each of the $m$ gaps in $c_{0}$, other chains would fill gaps. However, the structure of any possible packing of chains from $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ into a triple $(c_{i},c_{j},c_{k})$ would be a run of at least $a^{*}$ elements , followed potentially by a region of gaps with possible some elements in between (all of those gaps are size less than $a^{*}$), followed again by a run of at least $a^{*}$ elements, another region of gaps, and so on. This corresponds to each chain beginning with a consecutive sequence of at least $a^{*}$ elements. Because of this, it is not possible for the beginning of say chain $c_{j}$ to fill the gaps present in chain $c_{i}$. Therefore this alternate set of chains would not give way to a valid $13m+$ 1 solution, as it would be impossible to pack, and therefore there will not exist a bijection that satisfies Equation (9). Therefore by contradiction, the only possible $13m+$ 1 solution is to use the initially defined sets $\{c_{0},c_{1},...,c_{3m}\}$.
Therefore, if we have a $13m+$ 1 solution to Shor re-indexing, then it follows that we have successfully packed the elements of $A$ plus some constant $a^{*}$, encoded in unary, into $m$ gaps, each of size $T+3a^{*}$. From this point, we can recover a solution to the arbitray 3-partition problem instance.
@@
.4 Plots of Various Codes
We now present the compilation performance of both naive and Shor syndrome circuits of various error correction codes. Refer to Fig. 16 for the Steane code, and to Fig. 18 for the 3x3 toric code. These plots are similar to the style of Fig. 9, except we now have both the logical X and Z error rates corresponding to the top and bottom rows respectively. The left column corresponds to the performance on naive syndrome circuits, while the right column corresponds to performance on fault-tolerant Shor-syndrome circuits.
Figure 16: Compilation performance of the Steane code. The Steane code’s SSSC circuit is one of the few instances were assuming all gates commute did not work. However, the resulting SSSC circuit can be partitioned into 6 batches, one for each $\delta$ value. We fix this in the plot shown here by simply swapping two of those batches, allowing us to recover a functional syndrome extracting circuit.
Figure 17: Compilation performance of the Shor code.

### v2

@@ VI Optimal Compilation of regular LDPC codes
First we notice how a stabilizer tableau or parity check matrix maps to the input $\mathcal{S}$ of our ancillary qubit re-indexing problem. For example, take the Steane code (treating the X and Z checks together to create a more pedagogical example):
@@
We now notice that if our error correction code has regular column weight $w_{c}$ (each column has the same number of non-zero entries), then this would mean that all elements that occur in $\mathcal{S}$, occur the same number of times. Because each element occurs $w_{c}$ times, we can then make $w_{c}$ identical chains of the form n1,n2,,0n-1,n-2,...,0, each of which has no gaps, and will pack perfectly. Therefore, all LDPC codes with regular column weight can be compiled to run on $2\times n$ architectures in exactly $w_{c}$ shuttles. It is worth noting that this would then include all “regular” LDPC codes, as they have constant column and row weight [32]. Furthermore, the $[[144,12,12]]$ gross code  [25] also has regular column weight and would be an excellent choice of error correcting code for this architecture, as well as trivariate tricycle codes [26], generalized toric codes on twisted tori [29], lifted product codes and its’ subfamilies including generalized bicycle and two block group algebra codes [22]. In addition, we observed some irregular codes that also compiled perfectly such as $[[882,48,16]]$, and $[[882,12,?]]$ quasi-cyclic generalized hypergraph product codes [22].
VI.1 Optimal Compilation on Irregular LDPC codes
@@ VII Discussion and Conclusion
We have demonstrated methods for scheduling the shuttles of qubits in $2\times n$ one-dimensional architectures, highlighting a large family of LDPC codes that under Shor-syndrome extraction only ever requires a number of shuttles exactly equal to the code’s column weight, which notably includes bivariate bicycle quantum LDPC codes, which were shown in recent work to have a threshold on par with the surface code [25]. Other codes that would make great candidates for this architecture would include: trivariate tricycle codes [26], generalized toric codes on twisted tori [29], lifted product codes and its’ subfamilies including generalized bicycle and two block group algebra codes [22], as already mentioned in Section VI.
The most natural next problem to address would be the state preparation of “interwoven” cat states needed to perform Shor syndrome extraction. “Interwoven” corresponds to after ancilla reindexing, we might require the ancilla array contain distinct multiple cat states that need to be arranged in the ancilla array such that the physical qubits constituting one of the cat states are not all neighbors to one another (see Fig. 13). Throughout this work, we assumed that we had access to fault-tolerantly prepared cat states to load into the ancilla array, however, the details of how to do and optimize this, especially the process of interweaving, we leave to future work. In Appendix .3, we propose a sketch for how one could start with an array of consecutive cat states and interweave them, using an additional empty quantum dot array.
@@
For other future work, we plan to investigate the design of more elegant algorithms for compilation, perhaps exploring the trade-off between minimizing the size of the ancilla array and minimizing the number of shuttles. We will also consider other paradigms and pieces of the quantum error correction pipeline possible in the specific hardware implementation; for instance, incorporating flag qubits, architectures with more than two rails, how to optimize performance of logical operations, and discovery of other family codes that can have their syndrome extracted in ways that require only a few shuttling operations.
@@
Nevertheless, what we have presented here is already a substantial achievement towards embarking on what we believe to be a rich area of research, both from the perspectives of theoretical computer science as well as practical quantum error correction. That is, the rather novel algorithmic questions raised by optimizing error correcting codes to run on practical hardware are challenging enough to be interesting, yet seemingly constrained enough to allow for progress to be made. The culmination of which is the practical result of having more efficient quantum error correction schemes specifically tailored to a specific and promising type of quantum hardware. We hope to see this work expanded upon, as well as other works that explore these sorts of optimizations for other qubit platforms as we continue to approach a future of fault-tolerant quantum computation.
@@ Acknowledgements.
We would like to thank our colleagues Filip Rozepedek and Guus Avis for their fruitful discussions regarding gate commutativity. We also thank Nithin Raveendran for his recommendation to see how bivariate bicycle codes compile on this architecture. We would like to acknowledge colleagues Ferdinand Kuemmeth and Fabrizio Berritta as well. This work was supported by funding from the Inge Lehmann Programme of the Independent Research Fund Denmark, the US Army Research Office (ARO) under Award No. W911NF-24-2-0043, and the NSF under grants 1941583, 2346089, and 2402861.
@@
In this appendix, we provide examples of the block stacking visualization on valid error correcting codes, an illustration of how to translate between the “ladder” and “staircase” notations, a proof for the NP hardness of the ancillary qubit re-indexing problem, a discussion of how to prepare the necessary interwoven cat states needed for compiled Shor syndrome circuits, and a presentation of plotting logical versus physical error rates for different compilation schemes, following a circuit-level noise model.
@@
Refer to Figures 9 and 10 to see how the simple heuristics from Section III can be applied to the naive syndrome extracting circuits of Steane’s [[7,1,3]] code [18] and Shor’s [[9,1,3]] code [17]. In Section IV, we introduced a condensed version of the staircase notation that is specifically for Shor syndrome circuits. Fig. 11 illustrates how these two notations are equivalent.
Figure 9: Application of heuristics on the Steane code. Here we have the primitive sets of the naive syndrome extraction circuit of the Steane code expressed in block form. We can see that either of the first two heuristics discussed in Section III will allow use to find an optimal solution for this case of ancillary qubit re-indexing. Note: As this example is pedagogical in nature, this corresponds to reindexing the X and Z syndrome circuits together, something that in practice will be avoided.
Figure 10: Application of heuristics on the Shor code. Here we have the primitive sets of the naive syndrome extraction circuit of the Shor code expressed in block form. We can see that either of the first two heuristics discussed in Section III will allow use to find an optimal solution for this case of ancillary qubit re-indexing. Note: As this example is pedagogical in nature, this corresponds to reindexing the X and Z syndrome circuits together, something that in practice will be avoided.
@@
For this reduction, we take an arbitrary 3 partition problem and map it to our Shor re-indexing problem by mapping the elements of $A$ to a multiset $\mathcal{S}$, where $\mathcal{S}$ is the input to our problem as defined above. This $\mathcal{S}$ can then be decomposed into no fewer than $3m+1$ sets, each corresponding to a chain $c_{i}$, where $i\in\{0,1,...,3m\}$. Rather than specifying each element in $\mathcal{S}$, we instead define each $c_{i}$ below, noting that $\mathcal{S}=\sum_{i=0}^{3m}c_{i}$ (addition of multisets here is defined as combining them, e.g. $\{1,2\}+\{2,3\}=\{1,2,2,3\}$.) Later we prove that with how our $\mathcal{S}$ is defined, no other set of chains would be feasible in obtaining a solution to the reindexing problem, thus excluding the possibility of a solution to Shor re-indexing existing that does not also map to a valid solution of the given 3-partition problem. We define $c_{0},c_{1},...,c_{3m}$ as follows:
@@
Finally we notice that 0 occurs $3m$ times, so this construction needs at least $3m$ shuttles. Furthermore, due to 3) above, our construction will require a minimum of $3m+1$ shuttles, and our problem is to determine whether or not a bijection $\pi:[s]\rightarrow[s]$, recall $s=|\mathcal{S}|=\sum_{i=0}^{3m}|c_{i}|$, exists such that:
@@
To structure our proof, we first prove that if a solution exists in the 3-partition instance, then a solution of exactly $3m+1$ shuttles must exist for our problem. We prove afterwards that if a solution to our problem exists, then a solution to the 3-partition problem must also exist.
@@
Therefore each triple of chains would be packable into the gaps that occur within $c_{0}$. Therefore, all chains can be packed into $s$ positions exactly, implying the existence of a $3m+1$ shuttle solution.
@@
Therefore $m(T+4a^{*})-1$ is the number of spaces between the first and last element, where the 1-1 is due to the number of spaces between, for example, $2$ and 0 being $1$ in this context. As provided by Lemma 2, in those spaces, we have (m1)(m-1) ridges of size $a^{*}$, one ridge of size $a^{*}-1$. So there are a total of $(m-1)a^{*}+(a^{*}-1)$ elements in this spaces, meaning that the total size of all gaps in $c_{0}$ is given equal to :
@@
Direction 2: Shor Re-indexing $\rightarrow$ 3-Partition: Now we prove that any $3m+1$ shuttle solution to Shor re-indexing would also be a solution to the 3-partition problem. We prove this by proving that our construction could only ever provide a $3m+1$ shuttle solution if our definitions of $c_{0},c_{1},...,c_{3m}$ are followed exactly, which would then mean that each chain, aside from $c_{0}$, represents an element from $A$ encoded in unary, and as discussed in Direction 1, the packing of those chains corresponds to triples from $A$ each summing to $T$. In other words, we designed our reduction so that technically speaking, the idea of packing chains is not enforced, however by design, the only possible $3m+1$ shuttle solution will also be the solution to packing our chains as defined, and no other collection of chains would be feasible. To show that no other collection if feasible, we first prove the remaining two lemmas.
@@
We observe that $a^{*}<\frac{T}{4}$, and that $a<\frac{T}{2}$ for all $a\in A$. Therefore $\max_{e\in\sum_{i=1}^{3m}c_{i}}e<\frac{T}{4}+\frac{T}{2}$. The smallest value in $c_{0}$ is $|S|+T$, so if the largest element in $\sum_{i=1}^{3m}c_{i}$ and the smallest element in $c_{0}$ were in the same chain, a gap of size at least |S|+T/41|S|+T/4-1 would be created. Because we are looking for a bijection $\pi$ that maps from 1,,|S|1,...,|S| to 1,,|S|1,...,|S|, the longest possible chain (including its gaps) would be limited to length $|S|$, and the largest gap possible would be |S|2|S|-2. Therefore, because |S|+T/41>|S|2|S|+T/4-1>|S|-2, as $T/4>0$, it would not be possible for any element of $c_{o}$ to participate in one of the other chains. ∎
@@
We use the above lemmas to prove that no other collection of chains would result in a $3m+1$ solution. For the sake of contradiction, let’s assume that such an alternate set exists. By Lemma 3, we know that $c_{0}$ must remain unchanged, as $c_{0}$ constitutes $1$ shuttle, and the other $3m$ shuttles must come from the other elements, as 0 occurs $3m$ times. Therefore this alternate set of chains would be $\{c_{0},c_{1}^{\prime},...,c_{3m}^{\prime}\}$, and a gap must occur at least once in a chain in $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ by Lemma 4. By Lemma 1, we know there are $m$ gaps of size $T+3a^{*}$, and thus to have a $3m+1$ solution, we must be able to pack $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ into those $m$ gaps of size $T+3a^{*}$. Therefore the gaps introduced by using an alternate set of chains must be filled by other chains. As all gaps are by definition of size less than $a^{*}$, the ridges (Lemma 2) within $c_{0}$ cannot be used to fill them, which then leaves the only possibility to be that within each of the $m$ gaps in $c_{0}$, other chains would fill gaps. However, the structure of any possible packing of chains from $\{c_{1}^{\prime},...,c_{3m}^{\prime}\}$ into a triple $(c_{i},c_{j},c_{k})$ would be a run of at least $a^{*}$ elements , followed potentially by a region of gaps with possible some elements in between (all of those gaps are size less than $a^{*}$), followed again by a run of at least $a^{*}$ elements, another region of gaps, and so on. This corresponds to each chain beginning with a consecutive sequence of at least $a^{*}$ elements. Because of this, it is not possible for the beginning of say chain $c_{j}$ to fill the gaps present in chain $c_{i}$. Therefore this alternate set of chains would not give way to a valid $3m+1$ solution, as it would be impossible to pack, and therefore there will not exist a bijection that satisfies Equation (9). Therefore by contradiction, the only possible $3m+1$ solution is to use the initially defined sets $\{c_{0},c_{1},...,c_{3m}\}$.
Therefore, if we have a $3m+1$ solution to Shor re-indexing, then it follows that we have successfully packed the elements of $A$ plus some constant $a^{*}$, encoded in unary, into $m$ gaps, each of size $T+3a^{*}$. From this point, we can recover a solution to the arbitray 3-partition problem instance.
@@
.4 Simulation Results
Through simulation, we show that our methods of circuit compilation can be used to reduce the threshold of various error correcting codes. We model Pauli noise with highly efficient Pauli-Frame simulation techniques [34, 35] implemented in Julia [14]. We simulate at least 20,000 samples for each set of noise parameter values (each data point in our plots) in order to calculate the logical error rate for a single round of error correction. The simulated circuits contain the following steps:
  • Logical operator measurement In order to have something to compare our after-decoding results to, we measure the logical X or Z operators at the end of each syndrome circuit, as if it were an additional parity check. More specifically, we cannot measure both operators in the same circuit, so we need to run two simulations in order to obtain both the logical X and Z error rates.
[…]
Fig. 15 presents the compilation performance of both naive and Shor syndrome circuits for the 3x3 toric code. These plots are similar to the style of Fig. 14, except we now have both the logical X and Z error rates corresponding to the top and bottom rows of the figure respectively. The left column corresponds to the performance on naive syndrome circuits, while the right column corresponds to performance on fault-tolerant Shor-syndrome circuits.